OpenStudy (anonymous):

Y = X*X - 6X + 8 Let A = the value when X is Zero Let B = the largest value when Y = 0 Let C = the smaller value when Y = 0 THEN GO TO: N27 26.A(3*C)(A+1)W82 33.A(2*B)C+1

OpenStudy (anonymous):

what is this? N27 26.A(3*C)(A+1)W82 33.A(2*B)C+1

OpenStudy (anonymous):

Coordinates for a Geocache.

OpenStudy (anonymous):

I can figure out the coordinates equation but the first part I don't know how to do.

OpenStudy (anonymous):

I can help yo find A, B, and C. For A plug in 0 for x. For B and C, factor \(x^2-6x+8=0\) Find two numbers that multiply to 8 and add up to -6

OpenStudy (anonymous):

I understand B and C, but I'm still confused about A. Could you explain that further please?

OpenStudy (anonymous):

You substitute 0 for the x in the equation, then simplify to get y Y=0² -6(0)+8

OpenStudy (anonymous):

Y=8?

OpenStudy (anonymous):

yes. so A = 8 What did you get for B and C?

OpenStudy (anonymous):

B=-2 C=-4

OpenStudy (anonymous):

Start with this \(x^2-6x+8=0\) Use the two numbers you found to factor \((x-4)(x-2)=0\) Set each of the factors equal to 0 \(x-4=0\) or \(x-2=0\) Solve the equations for x \(x=4\), \(x=2\) So B = 4, C = 2

OpenStudy (anonymous):

Oh that makes sense thank you very much!

OpenStudy (anonymous):

you're welcome